Can you help me solve my word problem...instructions...the cost in dollars of manufacturing (w)wing nuts is given by the xpression 0.07w + 13.3. Find the cost when 375 wing nuts are made. What is the average cost to manufacture one wing nut?
Answer by bucky(2189)

The cost (C) in dollars of manufacturing a number of wing nuts (W) is 7 cents (0.07 dollars)
plus $13.30. In equation form this is:
.
C = 0.07*W + 13.30
.
The number of wing nuts, is 375. Substitute this number for W in the equation to get:
.
C = 0.07*375 + 13.30
.
Multiply out the first term on the right side and add it to 13.30 gives you:
.
C = 26.25 + 13.30 = $39.55
What is the average cost to make a wing nut? Depends. In this case, the manufacturing
cost is $39.55 for 375 wing nuts. By dividing 375 into $39.55 you find that each wing
nut costs about 10.55 cents to make. Suppose the customer only wanted 1 wing nut built. in that case
the equation is:
.
C = 0.07*1 + 13.30 = 13.37 for 1 nut or $13.37 each to make.
.
However, if the customer wanted a million wing nuts the total cost of manufacture
would be:
.
C = 0.07*1,000,000 + 13.30 = $70,013.30
.
And when you divide that by the quantity 1 million, you find that each wing nut costs
an average of 7.00133 cents to make.
.
So the question on the average cost for a wing nut really depends on the quantity
of wing nuts ordered. I guess that you have to assume that the question refers to the
customer's order of 375 wing nuts in which case the cost to manufacture is 10.55 cents for
each wing nut.